# MultiGPrompt for Multi-Task Pre-Training and Prompting on Graphs

### 14 Dec 2023 v2

@@ Abstract.
Graphs can inherently model interconnected objects on the Web, thereby facilitating a series of Web applications, such as web analyzing and content recommendation. Recently, Graph Neural Networks (GNNs) have emerged as a mainstream technique for graph representation learning. However, their efficacy within an end-to-end supervised framework is significantly tied to the availability of task-specific labels. To mitigate labeling costs and enhance robustness in few-shot settings, pre-training on self-supervised tasks has emerged as a promising method, while prompting has been proposed to further narrow the objective gap between pretext and downstream tasks. Although there has been some initial exploration of prompt-based learning on graphs, they primarily leverage a single pretext task, resulting in a limited subset of general knowledge that could be learned from the pre-training data. Hence, in this paper, we propose MultiGPrompt, a novel multi-task pre-training and prompting framework to exploit multiple pretext tasks for more comprehensive pre-trained knowledge. First, in pre-training, we design a set of pretext tokens to synergize multiple pretext tasks. Second, we propose a dual-prompt mechanism consisting of composed and open prompts to leverage task-specific and global pre-training knowledge, to guide downstream tasks in few-shot settings. Finally, we conduct extensive experiments on six public datasets to evaluate and analyze MultiGPrompt.
Graph neural networks, prompting, pre-training, multi-task, few-shot learning.
Part of the work was done while at Singapore Management University.
@@
ccs: Information systems Web miningccs: Information systems Data miningccs: Computing methodologies Learning latent representations
@@ 1. Introduction
The World Wide Web has evolved into an universal data repository, linking an expansive array of entities to create vast and intricate graphs. Mining such widespread graph data has fueled a myriad of Web applications, ranging from Web mining (Xu et al., 2022; Agarwal et al., 2022) and social network analysis (Zhang et al., 2022; Zhou et al., 2023) to content recommendation (Qu et al., 2023; Zhu et al., 2023). Contemporary techniques for graph analysis predominantly rely on graph representation learning, particularly graph neural networks (GNNs) (Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Xu et al., 2019). Most GNNs operate on a message-passing framework, where each node updates its representation by iteratively receiving and aggregating messages from its neighbors (Wu et al., 2020a), while more recent approaches have also explored transformer-based architectures (Yun et al., 2019; Hu et al., 2020a; Ying et al., 2021).
@@
Pre-training. GNNs are conventionally trained in an end-to-end manner, which heavily depends on the availability of large-scale, task-specific labeled data. To reduce the dependency on labeled data, there has been a growing emphasis on pre-training GNNs (Hu et al., 2020b, c; Qiu et al., 2020; Lu et al., 2021), following the success of pre-training in natural language and computer vision domains (Erhan et al., 2010; Dong et al., 2019; Bao et al., 2022). Pre-training GNNs involves optimizing self-supervised pretext tasks on graphs without task-specific labels. Such pretext tasks are designed to capture intrinsic graph properties, such as node and edge features (Hu et al., 2020c, b; You et al., 2020), node connectivity (Hamilton et al., 2017; Hu et al., 2020b; You et al., 2020; Lu et al., 2021), and local or global graph patterns (Veličković et al., 2018; You et al., 2020; Qiu et al., 2020; Hu et al., 2020c; Lu et al., 2021). Hence, pre-training yields a task-agnostic foundation that can be subsequently fine-tuned to a specific downstream task with a limited amount of labeled data. To further mitigate the inconsistency between pre-training and fine-tuning objectives (Liu et al., 2021b), prompt-based learning has been first proposed in language models (Brown et al., 2020). A prompt acts as an intermediary that reformulates the downstream task to align with pre-training, without the need to update the parameters of the pre-trained model. As a prompt entails much fewer parameters than the pre-trained model, it is especially amenable to few-shot settings where there are very limited task-specific labels.
@@
Following the success of prompt-based learning, researchers have also begun to explore prompt-based or related parameter-efficient learning on graph data (Liu et al., 2023a; Sun et al., 2023, 2022; Tan et al., 2023; Fang et al., 2022b; Li et al., 2023; Yu et al., 2024). However, most existing research in prompt-based graph learning only utilizes a single pretext task in pre-training. Not surprisingly, different pretext tasks capture different self-supervised signals from the graph. For example, link prediction tasks are more concerned with the connectivity or relationship between nodes (Sun et al., 2022), node/edge feature-based tasks focus more on the feature space (Tan et al., 2023), and subgraph-based tasks focus more on local or global information (Liu et al., 2023a; Yi et al., 2023). To cater to diverse downstream tasks, the pre-training step should aim to broadly extract knowledge from various aspects, such as node connectivity, node or edge features, and local or global graph patterns. Hence, it is ideal to incorporate multiple pretext tasks in pre-training in order to cover a comprehensive range of knowledge.
@@
First, in the pre-training stage, how can we leverage diverse pretext tasks for graph models in a synergistic manner? A straightforward way is to sum the losses of multiple pretext tasks directly, which have been explored in language (Devlin et al., 2018; Su et al., 2021; Wu et al., 2021), vision (Mormont et al., 2020; Gong et al., 2021) and graph data (Hu et al., 2019, 2020b; Lu et al., 2021; Fan et al., 2023). However, several works on multi-task learning (Wu et al., 2020b; Yu et al., 2020; Wang et al., 2022b) observe frequent task interference when tasks are highly diverse, resulting in suboptimal pre-training. A recent work (Wang et al., 2022b) sheds light on more synergistic multi-task pre-training in language models via an improved transformer design, which uses an attention mask and weighted skip connection to reduce task interference. Nonetheless, it largely remains an open problem for graph models.
Second, in the adaptation stage, how can we transfer both task-specific and global pre-trained knowledge to downstream tasks? Multiple pretext tasks further complicates the alignment of downstream objectives with the pre-trained model. On one hand, recent prompt-based learning on graphs (Liu et al., 2023a; Sun et al., 2022; Fang et al., 2022b) only focus on the downstream adaptation to a single pretext task. On the other hand, for language models a prompt-aware attention module has been incorporated into the transformer architecture (Wang et al., 2022b) to focus on extracting task-specific information from pre-training, lacking a global view of the pre-trained knowledge.
@@
Toward the first challenge, we draw inspiration from multi-task prompt-aware language models (Wang et al., 2022b), and design a series of pretext tokens to synergize the pretext tasks during pre-training. As illustrated in Fig. 1(a), we associate each pretext task with one (or more) task-specific pretext tokens, which are used to reformulate the input of the pretext task. A pretext token is simply a learnable vector, and thus yields a learnable reformulation in a task-specific fashion. The reformulation guides multiple pretext tasks into a synergistic integration that enables collaboration, rather than interference, among the diverse tasks.
Towards the second challenge, we propose a dual-prompt mechanism, employing both a composed prompt and an open prompt to harness task-specific and global pre-training knowledge, respectively, as shown in Fig. 1(b). More specifically, a composed prompt is a learnable composition (e.g., a linear or neural network-based combination) of the pre-trained (frozen) pretext tokens, similar to the approach in language models (Wang et al., 2022b). As the composed prompt builds upon the pretext tokens, it is designed to query the pre-trained model for a precise mixture of information specific to each pretext task, focusing on task-specific pre-trained knowledge. However, it falls short of a global view to extract relevant inter-task knowledge (e.g., the relations or interactions between pretext tasks) from the whole pre-trained model. Hence, we propose the concept of open prompt, which aims to transfer global inter-task knowledge to complement the task-specific knowledge of the composed prompt.
@@ 2. Related Work
Graph pre-training. Borrowing insights from the realm of pre-training methodologies in both the language (Dong et al., 2019; Brown et al., 2020; Gao et al., 2020; Schick and Schütze, 2020) and vision (Bao et al., 2021; Zhou et al., 2022b, a; Zang et al., 2022) domains, a myriad of GNN-based pre-training approaches have emerged (Hu et al., 2019, 2020b; Lu et al., 2021; Qiu et al., 2020; Kipf and Welling, 2016). These methods leverage the intrinsic graph structures in a self-supervised manner, setting the stage for knowledge transfer to downstream tasks. This transfer can be accomplished by a fine-tuning process that capitalizes on labeled data pertinent to each downstream task.
@@
However, a gap emerges between the objectives of pre-training and fine-tuning (Liu et al., 2023b). On one hand, pre-training seeks to distill general knowledge from the graph without relying on explicit supervision. Conversely, fine-tuning tailors to specific supervisory signals aligned with the downstream tasks. This gap in objectives can hinder the transfer of knowledge from the pre-trained model, potentially hurting the downstream performance.
@@
Graph prompt learning. Originated in the language domain (Brown et al., 2020; Wei et al., 2021), prompt-based learning has been effective in bridging the gap between pre-training and downstream objectives. Specifically, prompts can be tuned for each downstream task, steering each task toward the pre-trained model while keeping the pre-trained parameters frozen. Due to the parameter-efficient nature of prompt, it has been quickly popularized in favor of fine-tuning larger pre-trained models, or when the downstream task only has few-shot labels. Given the advantages, prompt-based learning has also been explored on visual data (Jia et al., 2021; Zhou et al., 2022b; Jia et al., 2022) and graphs (Liu et al., 2023a; Sun et al., 2022; Fang et al., 2022b; Tan et al., 2023; Sun et al., 2023).
@@
Specific to graph data, GraphPrompt (Liu et al., 2023a) and ProG (Sun et al., 2023) attempt to unify pre-training and typical downstream tasks on graphs into a common template, and further introduce learnable prompts to guide the downstream tasks. Their difference mainly lies in learning scenarios: ProG requires a set of base classes in a typical meta-learning setup (Zhou et al., 2019; Wang et al., 2020; Liu et al., 2021a), while GraphPrompt does not make use of base classes. In contrast, GPPT (Sun et al., 2022) and VNT (Tan et al., 2023) only focus on the node classification task downstream. However, all these methods only employ a single pretext task, and thus lack a comprehensive coverage of pre-trained knowledge in different aspects.
@@
Multi-task pre-training. To broaden and diversify beyond a single pretext task, multi-task pre-training methods have been proposed for language (Devlin et al., 2018; Su et al., 2021; Wu et al., 2021), vision (Mormont et al., 2020; Gong et al., 2021) and graph data (Hu et al., 2019, 2020b; Lu et al., 2021; Fan et al., 2023). However, on one hand, these methods directly aggregate multiple losses from diverse pretext tasks, resulting in task interference in the pre-training stage (Wu et al., 2020b; Yu et al., 2020; Wang et al., 2022b). On the other hand, these approaches only perform fine-tuning for the downstream tasks in the adaptation stage, which is inadequate to align the multiple pretext tasks with the downstream objective. To mitigate these issues, a recent study (Wang et al., 2022b) employs prompts to integrate multiple pretext tasks and further guide downstream tasks. However, it is designed for language models, requiring specific modification to the transfomer architecture. Moreover, it lacks a global view over the multiple pretext tasks.
On another line, there is some research on multi-modal prompts (Khattak et al., 2023; Dong et al., 2024), employing multiple prompts to different modality of data such as vision and language. They aim to align the representations from different modalities, which diverges from our multi-task objectives in pre-training.
@@ 3. Preliminaries
Graph. A graph is represented as $G=(V,E)$, with $V$ denoting the set of nodes and $E$ the set of edges. Equivalently, the graph can be represented by an adjacency matrix $\mathbf{A}$, such as $\mathbf{A}_{ij}=1$ iff $(v_{i},v_{j})\in E$, for any $v_{i},v_{j}\in V$. We further consider an input feature matrix for the nodes, given by $\mathbf{X}\in\mathbb{R}^{|V|\times d}$. For a node $v_{i}\in V$, its feature vector is represented as $\mathbf{x}_{i}\in\mathbb{R}^{d}$.
For a dataset with multiple graphs, we use the notation $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$.
Graph encoder. GNNs are popular choices of graph encoder, most of which employ a message-passing mechanism (Wu et al., 2020a). Specifically, each node in the graph aggregates messages (i.e., input features or embeddings) from its neighboring nodes to update its own embedding. Multiple layers of neighborhood aggregation can be stacked, facilitating recursive message passing across the graph. Formally, let $\mathbf{H}^{l}$ be the embedding matrix of the graph at the $l$-th layer, where its $i$-th row, $\mathbf{h}_{i}^{l}$, corresponds to the embedding of node $v_{i}$. It is computed based on the embeddings from the preceding layer:
@@ 4.1. Overall Framework
Next, as shown in Fig. 2(b), we aim to transfer the pre-trained knowledge to different downstream tasks. We propose a dual-prompt mechanism with a series of composed prompts, $\mathcal{P}_{\langle\text{com}\rangle}$, and open prompts, $\mathcal{P}_{\langle\text{op}\rangle}$. The composed prompt is obtained by a learnable aggregation of the pretext tokens to leverage task-specific pre-trained knowledge. The open prompt, on the other hand, is a learnable vector that learns global inter-task insights from the pre-trained model. Both are then used to reformulate the downstream input to the pre-trained model separately, and their respective output embeddings are combined and further fed into the downstream task loss.

### 22 Feb 2024 v6

@@ Abstract.
Graph Neural Networks (GNNs) have emerged as a mainstream technique for graph representation learning. However, their efficacy within an end-to-end supervised framework is significantly tied to the availability of task-specific labels. To mitigate labeling costs and enhance robustness in few-shot settings, pre-training on self-supervised tasks has emerged as a promising method, while prompting has been proposed to further narrow the objective gap between pretext and downstream tasks. Although there has been some initial exploration of prompt-based learning on graphs, they primarily leverage a single pretext task, resulting in a limited subset of general knowledge that could be learned from the pre-training data. Hence, in this paper, we propose MultiGPrompt, a novel multi-task pre-training and prompting framework to exploit multiple pretext tasks for more comprehensive pre-trained knowledge. First, in pre-training, we design a set of pretext tokens to synergize multiple pretext tasks. Second, we propose a dual-prompt mechanism consisting of composed and open prompts to leverage task-specific and global pre-training knowledge, to guide downstream tasks in few-shot settings. Finally, we conduct extensive experiments on six public datasets to evaluate and analyze MultiGPrompt.
Graph learning, prompting, multi-task, few-shot learning.
$*{}^{*}$art of the work was done while at Singapore Management University.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589334.3645423isbn: 979-8-4007-0171-9/24/05ccs: Information systems Web miningccs: Information systems Data miningccs: Computing methodologies Learning latent representations
@@ 1. Introduction
The Web has evolved into an universal data repository, linking an expansive array of entities to create vast and intricate graphs. Mining such widespread graph data has fueled a myriad of Web applications, ranging from Web mining (Xu et al., 2022; Agarwal et al., 2022) and social network analysis (Zhang et al., 2022; Zhou et al., 2023) to content recommendation (Qu et al., 2023; Zhu et al., 2023). Contemporary techniques for graph analysis predominantly rely on graph representation learning, particularly graph neural networks (GNNs) (Kipf and Welling, 2017; Veličković et al., 2018; Fang et al., 2023a, b; Yu et al., 2023a). Most GNNs operate on a message-passing framework, where each node updates its representation by iteratively receiving and aggregating messages from its neighbors (Wang et al., 2023a, 2022b; Anonymous, 2024b, a; Wang et al., 2023b; Li et al., 2023, 2024b), while more recent approaches have also explored transformer-based architectures (Yun et al., 2019; Hu et al., 2020a; Wu et al., 2024).
@@
Pre-training. GNNs are conventionally trained in an end-to-end manner, which heavily depends on the availability of large-scale, task-specific labeled data. To reduce the dependency on labeled data, there has been a growing emphasis on pre-training GNNs (Hu et al., 2020b, c; Qiu et al., 2020; Lu et al., 2021; Yu et al., 2024a), following the success of pre-training in natural language and computer vision domains (Erhan et al., 2010; Dong et al., 2019; Bao et al., 2022). Pre-training GNNs involves optimizing self-supervised pretext tasks on graphs without task-specific labels. Such pretext tasks are designed to capture intrinsic graph properties, such as node and edge features (Hu et al., 2020c, b; You et al., 2020), node connectivity (Hamilton et al., 2017; Hu et al., 2020b; You et al., 2020; Lu et al., 2021), and local or global graph patterns (Veličković et al., 2018; You et al., 2020; Qiu et al., 2020; Hu et al., 2020c; Lu et al., 2021). Hence, pre-training yields a task-agnostic foundation that can be subsequently fine-tuned to a specific downstream task with a limited amount of labeled data. To further mitigate the inconsistency between pre-training and fine-tuning objectives (Liu et al., 2021b), prompt-based learning has been first proposed in language models (Brown et al., 2020). A prompt acts as an intermediary that reformulates the downstream task to align with pre-training, without the need to update the parameters of the pre-trained model. As a prompt entails much fewer parameters than the pre-trained model, it is especially amenable to few-shot settings where there are very limited task-specific labels.
@@
Following the success of prompt-based learning, researchers have also begun to explore prompt-based or related parameter-efficient learning on graph data (Liu et al., 2023a; Sun et al., 2023, 2022; Tan et al., 2023; Fang et al., 2022; Li et al., 2024a; Yu et al., 2024b). However, most existing research in prompt-based graph learning only utilizes a single pretext task in pre-training. Not surprisingly, different pretext tasks capture different self-supervised signals from the graph. For example, link prediction tasks are more concerned with the connectivity or relationship between nodes (Sun et al., 2022), node/edge feature-based tasks focus more on the feature space (Tan et al., 2023), and subgraph-based tasks focus more on local or global information (Liu et al., 2023a; Yi et al., 2023). To cater to diverse downstream tasks, the pre-training step should aim to broadly extract knowledge from various aspects, such as node connectivity, node or edge features, and local or global graph patterns. Hence, it is ideal to incorporate multiple pretext tasks in pre-training in order to cover a comprehensive range of knowledge.
@@
First, in the pre-training stage, how can we leverage diverse pretext tasks for graph models in a synergistic manner? A straightforward way is to sum the losses of multiple pretext tasks directly, which have been explored in language (Devlin et al., 2018; Su et al., 2021; Wu et al., 2021), vision (Mormont et al., 2020; Gong et al., 2021) and graph data (Hu et al., 2019, 2020b; Lu et al., 2021; Fan et al., 2023). However, several works on multi-task learning (Wu et al., 2020b; Yu et al., 2020; Wang et al., 2022c) observe frequent task interference when tasks are highly diverse, resulting in suboptimal pre-training. A recent work (Wang et al., 2022c) sheds light on more synergistic multi-task pre-training in language models via an improved transformer design, using an attention mask and weighted skip connection to reduce task interference. It largely remains an open problem for graph models.
Second, in the adaptation stage, how can we transfer both task-specific and global pre-trained knowledge to downstream tasks? Multiple pretext tasks further complicates the alignment of downstream objectives with the pre-trained model. On one hand, recent prompt-based learning on graphs (Liu et al., 2023a; Sun et al., 2022; Fang et al., 2022) only focus on the downstream adaptation to a single pretext task. On the other hand, for language models a prompt-aware attention module has been incorporated into the transformer architecture (Wang et al., 2022c) to focus on extracting task-specific information from pre-training, lacking a global view of the pre-trained knowledge.
@@
Toward the first challenge, we draw inspiration from multi-task prompt-aware language models (Wang et al., 2022c), and design a series of pretext tokens to synergize the pretext tasks during pre-training. As illustrated in Fig. 1(a), we associate each pretext task with one (or more) task-specific pretext tokens, which are used to reformulate the input of the pretext task. A pretext token is simply a learnable vector, and thus yields a learnable reformulation in a task-specific fashion. The reformulation guides multiple pretext tasks into a synergistic integration that enables collaboration, rather than interference, among the diverse tasks.
Towards the second challenge, we propose a dual-prompt mechanism, employing both a composed prompt and an open prompt to harness task-specific and global pre-training knowledge, respectively, as shown in Fig. 1(b). More specifically, a composed prompt is a learnable composition (e.g., a linear or neural network-based combination) of the pre-trained (frozen) pretext tokens, similar to the approach in language models (Wang et al., 2022c). As the composed prompt builds upon the pretext tokens, it is designed to query the pre-trained model for a precise mixture of information specific to each pretext task, focusing on task-specific pre-trained knowledge. However, it falls short of a global view to extract relevant inter-task knowledge (e.g., the relations or interactions between pretext tasks) from the whole pre-trained model. Hence, we propose the concept of open prompt, which aims to transfer global inter-task knowledge to complement the task-specific knowledge of the composed prompt.
@@ 2. Related Work
Graph pre-training. GNN-based pre-training approaches (Hu et al., 2019, 2020b) leverage the intrinsic graph structures in a self-supervised manner, setting the stage for knowledge transfer to downstream tasks. This transfer can be achieved by a fine-tuning process that capitalizes on labeled data pertinent to each downstream task.
@@
However, a gap emerges between the objectives of pre-training and fine-tuning (Liu et al., 2023b). On one hand, pre-training seeks to distill general knowledge from the graph without supervision. Conversely, fine-tuning tailors to specific supervisory signals aligned with the downstream tasks. This gap in objectives can hinder the transfer of knowledge, potentially hurting the downstream performance.
@@
Graph prompt learning. Prompt-based learning has been effective in bridging the gap between pre-training and downstream objectives. Specifically, prompts can be tuned for each downstream task, steering each task toward the pre-trained model while keeping the pre-trained parameters frozen. Due to the parameter-efficient nature of prompt, it has been quickly popularized in favor of fine-tuning larger pre-trained models, or when the downstream task only has few-shot labels. Given the advantages, prompt-based learning has also been explored on graphs (Yu et al., 2024a; Liu et al., 2023a; Sun et al., 2022; Tan et al., 2023; Sun et al., 2023).
@@
Specifically, GraphPrompt (Liu et al., 2023a) and ProG (Sun et al., 2023) attempt to unify pre-training and typical downstream tasks on graphs into a common template, and further introduce learnable prompts to guide the downstream tasks. Note that ProG requires a set of base classes in a typical meta-learning setup (Zhou et al., 2019; Wang et al., 2020), while GraphPrompt does not make use of base classes. In contrast, GPPT (Sun et al., 2022) and VNT (Tan et al., 2023) only focus on the node-level task. However, all these methods only employ a single pretext task, and thus lack a comprehensive coverage of pre-trained knowledge in different aspects.
@@
Multi-task pre-training. To broaden and diversify beyond a single pretext task, multi-task pre-training methods have been proposed for graph data (Lu et al., 2021; Fan et al., 2023). However, on one hand, these methods directly aggregate multiple losses from diverse pretext tasks, resulting in task interference in the pre-training stage (Wu et al., 2020b; Wang et al., 2022c). On the other hand, they only perform fine-tuning for the downstream tasks in the adaptation stage, which is inadequate to align the multiple pretext tasks with the downstream objective. To mitigate these issues, a recent study (Wang et al., 2022c) employs prompts to integrate multiple pretext tasks and further guide downstream tasks. However, it is designed for language models, requiring specific modification to the transfomer architecture. Moreover, it lacks a global view over the multiple pretext tasks. Besides, there is a research on multi-modal prompts (Khattak et al., 2023; Dong et al., 2024), employing multiple prompts to different modality of data such as vision and language. It aims to align the representations from different modalities, which diverges from our multi-task objectives in pre-training.
@@ 3. Preliminaries
Graph. A graph is represented as $G=(V,E)$, with $V$ denoting the set of nodes and $E$ the set of edges. Equivalently, the graph can be represented by an adjacency matrix $\mathbf{A}$, such as $\mathbf{A}_{ij}=1$ iff $(v_{i},v_{j})\in E$, for any $v_{i},v_{j}\in V$. We further consider an input feature matrix for the nodes, given by $\mathbf{X}\in\mathbb{R}^{|V|\times d}$. For a node $v_{i}\in V$, its feature vector is represented as $\mathbf{x}_{i}\in\mathbb{R}^{d}$. For a dataset with multiple graphs, we use the notation $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$.
Graph encoder. GNNs are popular choices of graph encoder, most of which employ a message-passing mechanism (Wu et al., 2020a). Specifically, each node in the graph aggregates messages from its neighboring nodes to update its own embedding. Multiple layers of neighborhood aggregation can be stacked, facilitating recursive message passing across the graph. Formally, let $\mathbf{H}^{l}$ be the embedding matrix of the graph at the $l$-th layer, where its $i$-th row, $\mathbf{h}_{i}^{l}$, corresponds to the embedding of node $v_{i}$. It is computed based on the embeddings from the preceding layer:
@@ 4.1. Overall Framework
Next, as shown in Fig. 2(b), we aim to transfer the pre-trained knowledge to different downstream tasks. We propose a dual-prompt mechanism with a series of composed prompts, $\mathcal{P}_{\langle\text{com}\rangle}$, and open prompts, $\mathcal{P}_{\langle\text{op}\rangle}$. The composed prompt is obtained by a learnable aggregation of the pretext tokens to leverage task-specific pre-trained knowledge. The open prompt is a learnable vector that learns global inter-task insights from the pre-trained model. Both are then used to reformulate the downstream input to the pre-trained model separately, and their respective output embeddings are combined and further fed into the downstream task loss.